# Knowledge Distillation Driven Semantic NOMA for Image Transmission with Diffusion Model

### v1

@@ Abstract
Index Terms:
Semantic communication, non-orthogonal multiple access, channel adaptation, knowledge distillation, diffusion model.
@@ I Introduction
With the gradual advancement of 6G wireless networks, wireless communication is entering a transformative era of technological innovation. Future 6G networks will not only accommodate an explosive growth in data traffic but also meet escalating user demands for intelligent and efficient connection services [1]. In this context, traditional bit-oriented communication paradigm has gradually revealed limitations. Among the promising candidates for 6G, semantic communication has garnered significant attention due to its potential to transcend the constraints of conventional bit-level transmission [2]. By integrating artificial intelligence, semantic communication shifts transmission objectives from bit-level to semantic-level information delivery, demonstrating superior transmission efficiency over traditional separated source-channel coding methods in bandwidth-constrained and complex channel environments. However, designing end-to-end semantic communication systems for multi-user scenarios face significant challenges: adapting to diverse channel conditions (e.g., additive white Gaussian noise [AWGN] channel and Rayleigh fading channel), mitigating non-orthogonal feature interference, and enhancing perceptual quality under strict bandwidth constraints [3]. Specifically, non-orthogonal transmission, where multiple users’ semantic features share the same time-frequency resources, significantly exacerbates interference, while diverse channels and low compression ratios degrade signal reconstruction quality. This paper presents a novel solution to address these challenges, thereby advancing the practical deployment of semantic communication in 6G networks.
I-A Related Works
I-A1 DeepJSCC in Single-User and Multi-User Semantic Communication
@@
Semantic communication has achieved remarkable progress in single-user scenarios, with numerous studies employing deep joint source-channel coding (DeepJSCC) to enable efficient semantic transmission of diverse modalities [4, 5]. The paper [6] introduces attention modules to adaptive signal-to-noise ratio (SNR) in DeepJSCC codec, and the work [7] introduces learnable entropy models to implement joint source-channel coding (JSCC) with variable-length transmission mechanism. These works underscore the potential of semantic communication in terms of dynamic performance and robustness. However, research on multi-user semantic communication remains in its infancy, with design complexity far exceeding that of single-user systems. In multi-user settings, semantic interference, limited transmission resources, and inter-user channel contention emerge as primary bottlenecks. The authors of [8] propose task-oriented frameworks for multi-user semantic communication, while the authors of [9] explores the feasibility of multimodal semantic fusion. The paper [10] combines beamforming algorithm and semantic coding in the downlink massive multiple-input multiple-output (MIMO) scenario to achieve multi-user image downlink semantic transmission. Nonetheless, these approaches predominantly rely on orthogonal resource allocation, resulting in a sharp increase in physical-layer overhead as the number of users grows. To address this, non-orthogonal multiple access (NOMA) concepts have been introduced. For instance, the paper [11] proposes the end-to-end (E2E) uplink DeepJSCC-NOMA network for wireless image transmission; the paper [12] leverages NOMA with quantization modules to enable intelligent multi-user detection; the work [13] employs semantic differential superposition coding at the transmitter and interference cancellation at the receiver to support both homogeneous and heterogeneous transmission (i.e., semantic information and bit streams); the paper [14] extracts and decorrelates basic and enhanced source features, then superposes them probabilistically for transmission, allowing receivers to decode semantic information commensurate with channel quality. However, these methods are still insufficient in terms of interference optimization in NOMA scenarios and cannot be applied to dynamic channel conditions. Additionally, the authors of [15] proposes Swin-Transformer-based feature fusion schemes under degraded broadcast channels (DBC) to reduce transmission overhead via joint representations, though the high computational complexity limits scalability in multi-user scenarios. Novel semantic multiple-access schemes have also emerged: the paper [16] exploits DeepJSCC’s interference resilience to propose an orthogonal model division approach, while the work [17] introduces semantic feature multiple access for multi-user video transmission. Despite integrating traditional multiple-access principles with neural network robustness, these solutions lack systematic strategies for interference elimination and resource optimization in multi-user non-orthogonal settings. This paper focuses on uplink multi-user non-orthogonal semantic transmission and designs a low-complexity deep learning framework that can adapt to different channel conditions and further optimize to reduce the impact of interference.
I-A2 Knowledge Distillation and Its Application in Semantic Communication
I-A3 Diffusion Model and Its Application in Semantic Communication
Conventional DeepJSCC frameworks typically aim to minimize pixel-level distortion, such as mean squared error (MSE), as their optimization objective, but they ignore the improvement of image perceptual quality, especially in poor channel environments. In recent years, generative models have achieved groundbreaking progress in computer vision, with diffusion models emerging as a superior alternative to GANs due to their capacity to produce high-quality and semantically coherent images. Within the domain of semantic communication, GANs have been employed to improve image transmission quality. For instance, the paper [24] models the entire DeepJSCC transmission pipeline as the forward process, leveraging a pretrained StyleGAN2 to solve the inverse problem and improve perceptual quality. In contrast, diffusion models offer distinct advantages in stability, resistance to mode collapse, and generative diversity. Accordingly, the authors of [25] introduce the diffusion-aided JSCC framework, utilizing a pretrained Stable Diffusion model to the image reconstruction with initial JSCC decoding and multimodal conditions (spatial, textual, and channel state information). Similarly, the works of [26] and [27] deploy pretrained diffusion models at the receiver to boost perceptual quality, Furthermore, the work [28] proposes channel denoising diffusion models (CDDM), which leverage diffusion models to preprocess channel outputs. The paper [29] presents the semantics-guided diffusion DeepJSCC (SGD-JSCC) scheme, which integrates diffusion models with multi-modal semantic guidance (i.e., text and edge maps) to design a channel denoising model. Inspired by these advancements, this paper innovatively proposes a two-stage image reconstruction framework tailored for semantic NOMA scenarios, deploying pretrained diffusion models at the receiver to achieve superior perceptual quality.
I-B Motivation and Contributions
@@
Although prior studies have leveraged deep learning techniques to design multi-user semantic communication systems, significant challenges persist in the context of uplink semantic NOMA transmission scenarios. These include adapting to dynamic channel conditions such as AWGN and Rayleigh fading channels, inadequate management of non-orthogonal semantic feature interference, and limited enhancements in multi-user perceptual image quality under strict bandwidth constraints. Dynamic channels degrade transmission performance, interference from superimposed different users’ signals reduces pixel-level fidelity, and high compression ratios and poor channel conditions impair perceptual metrics. These limitations hinder scalable 6G network deployment. Addressing these challenges in semantic NOMA communication, this paper proposes a novel framework that enhances image recovery performance across both pixel-level fidelity and perceptual quality. The main contributions of this work are summarized as follows:
@@
The rest of this paper is organized as follows. In Section II, we introduce the system model and describe the overall structure of SemNOMA framework. In Section III, we introduce the knowledge distillation strategy to optimize SemNOMA. In Section IV, we show that the diffusion prior can further recover the details of the reconstructed images. Numerical simulation results are provided in Section V, followed by the conclusion of this work in Section VI.
@@
II System Model
II-A Overview
@@
The proposed KDD-SemNOMA framework is illustrated in Figure 1. Consider an uplink multi-user image semantic communication scenario where multiple user equipment (UE) transmit uplink image semantic information to the base station (BS). Unlike conventional communication systems employing separate source and channel coding, our framework adopts JSCC to directly transmit semantic features. Both UE and BS are configured with single antennas. Each UE transmits an image $\mathbf{x}_{i}\in\mathbb{R}^{C\times H\times W}$, where $i$ denotes the index of UE, $C$, $H$, and $W$ represent the color channels, width, and height of the RGB image, respectively. We denote $m=C\times H\times W$. The transmitter encodes the source image $\mathbf{x}_{i}$ into compressed semantic features $\mathbf{s}_{i}\in\mathbb{C}^{k}$ through a JSCC encoder, which can be expressed as
@@
Our framework for multi-user image reconstruction at the receiver comprises two primary stages: (1) multi-user DeepJSCC decoding optimized by knowledge distillation as detailed in Section III, and (2) image refinement guided by diffusion-based generative priors as detailed in Section IV. At the first stage, the SemNOMA model is trained using knowledge distillation to effectively decode transmitted signals. At the second stage, we leverage pretrained diffusion model priors to refine low-quality received images, where their visual fidelity and detail have further enhanced.
II-B Enhanced AF-Module
@@
Let the input semantic feature map be $\mathbf{f}_{s}\in\mathbb{R}^{B\times C^{\prime}\times H^{\prime}\times W^{\prime}}$, where $B$, $C^{\prime}$, $H^{\prime}$, and $W^{\prime}$ denote batch size, number of channels, height, and width, respectively. The channel side information is represented by $\mathbf{s}\in\mathbb{R}^{B\times D}$, with $D=3$ corresponding to the SNR $\gamma\in[0,20]$ dB and Rayleigh fading parameters of the $i$-th UE (amplitude $a_{i}$ and phase $\phi_{i}$). During training, $\gamma$ is uniformly sampled from $[0,20]$ dB, and the Rayleigh parameters $(a_{i},\phi_{i})$ are randomly generated to simulate diverse channel conditions. This enhances the model’s adaptability and reduces the impact of channel on the semantic features.
@@
II-C SemNOMA Network Architecture
@@
The encoder processes RGB images and normalizes them to the $[0,1]$ range by dividing pixel values by 255. It employs a stem layer for initial downsampling, followed by LayerNorm and additional downsampling stages. Multiple ConvNeXt blocks, with $N_{e}$ blocks per stage, process features, as detailed in Table I for $\times32$ resolution and Table II for $\times256$ resolution. Enhanced AF-Module, as introduced in Section II-B, improve the robustness to channel noise by conditioning on SNR and Rayleigh fading parameters (i.e., $\gamma,a,\phi$). The final $1\times 1$ convolution produces $M$ channels, normalized to meet the power constraint (i.e., Eq. 2), and reshaped into $k$ complex-valued channel inputs.
The decoder reconstructs image from the received multi-user superimposed semantic features. The $k$ complex-valued samples are converted to $2k$ real-valued channels, processed by a $1\times 1$ convolution and LayerNorm. Progressive upsampling stages then restore spatial resolution (e.g., $32\times 32$ or $256\times 256$), each followed by $N_{d}$ ConvNeXt blocks with parametric rectified linear unit (PReLU) activations and AF-Modules for feature refinement, as specified in Table I and Table II. A final $1\times 1$ convolution ($K=3$) with sigmoid activation produces a $C=3$ output, which is then denormalized to range $[0,255]$.
@@
III Knowledge Distillation Optimization for Semantic NOMA (KD-SemNOMA)
@@
1:Multi-user image dataset $\mathbf{x}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, hyperparameters $\lambda_{1},\lambda_{2},\lambda_{3}$
@@
11:for each batch $\mathbf{x}_{\text{batch}}\subseteq\mathbf{x}$ do
@@
III-A Teacher Model Training

### v2

@@ I Introduction
With the gradual advancement of 6G wireless networks, wireless communication is entering a transformative era of technological innovation. Future 6G networks will not only accommodate an explosive growth in data traffic but also meet escalating user demands for intelligent and efficient connection services [1, 2, 3]. In this context, traditional bit-oriented communication paradigm has gradually revealed limitations[4]. Among the promising candidates for 6G, semantic communication has garnered significant attention due to its potential to transcend the constraints of conventional bit-level transmission [5]. By integrating artificial intelligence, semantic communication shifts transmission objectives from bit-level to semantic-level information delivery, demonstrating superior transmission efficiency over traditional separated source-channel coding methods in bandwidth-constrained and complex channel environments [6].
Semantic communication has achieved remarkable progress in single-user scenarios. On the one hand, numerous studies have adopted end-to-end deep joint source-channel coding (DeepJSCC) frameworks to enable efficient semantic transmission of diverse modalities [6, 7]. On the other hand, another line of research has integrated generative models to enhance semantic reconstruction quality, particularly under challenging channel conditions [8, 9, 10, 11].The paper [10] proposes a generative video semantic communication framework, the work [11] incorporates adversarial loss into a DeepJSCC-based orthogonal frequency division multiplexing (OFDM) system to enhance the perceptual quality of reconstructed images. These generative approaches have demonstrated superior performance in perceptual metrics, effectively compensating for the limitations of traditional distortion-based measures. Further improvements in dynamic adaptability and robustness are investigated in [12], which introduces attention modules for signal-to-noise ratio (SNR) adaptation in DeepJSCC codecs, and [13], which employs learnable entropy models for variable-length joint source-channel coding (JSCC).
@@
In contrast, research on multi-user semantic communication remains in its infancy, with design complexity far exceeding that of single-user systems. In multi-user settings, semantic interference, limited transmission resources, and inter-user channel contention emerge as primary bottlenecks. The authors of [14] propose task-oriented frameworks for multi-user semantic communication, while the authors of [15] explores the feasibility of multimodal semantic fusion. The paper [16] combines beamforming algorithm and semantic coding in the downlink massive multiple-input multiple-output (MIMO) scenario to achieve multi-user image downlink semantic transmission. Nonetheless, these approaches predominantly rely on orthogonal resource allocation, resulting in a sharp increase in physical-layer overhead as the number of users grows. To address this, non-orthogonal multiple access (NOMA) concepts have been introduced. For instance, the paper [17] proposes the end-to-end (E2E) uplink DeepJSCC-NOMA network for wireless image transmission. These works [18, 19, 20, 21] also design NOMA based semantic communication system with some interference management optimization methods. However, these methods are still insufficient in terms of interference mitigation in NOMA scenarios.
I-A Motivation and Contributions
@@
Although prior studies have leveraged deep learning techniques to design multi-user semantic communication systems, significant challenges persist in the context of uplink semantic NOMA transmission scenarios. These include adapting to dynamic channel conditions such as additive white Gaussian noise (AWGN) and Rayleigh fading channels, inadequate management of non-orthogonal semantic feature interference, and limited enhancements in multi-user perceptual image quality under strict bandwidth constraints. Dynamic channels degrade transmission performance, interference from superimposed different users’ signals reduces pixel-level fidelity, and high compression ratios and poor channel conditions impair perceptual metrics. These limitations hinder scalable 6G network deployment. Addressing these challenges in semantic NOMA communication, this paper proposes a novel framework that enhances image recovery performance across both pixel-level fidelity and perceptual quality. The main contributions of this work are summarized as follows:
@@
The rest of this paper is organized as follows. In Section III, we introduce the system model and describe the overall structure of SemNOMA framework. In Section IV, we introduce the knowledge distillation strategy to optimize SemNOMA. In Section V, we show that the diffusion prior can further recover the details of the reconstructed images. Numerical simulation results are provided in Section VI, followed by the conclusion of this work in Section VII.
@@
II Related Works
This section reviews recent advances in semantic communication, focusing on three key areas: multi-user transmission techniques, knowledge distillation applications in semantic communication, and diffusion models for semantic enhancement.
II-A DeepJSCC in Multi-User Semantic Communication
In multi-user settings, semantic interference, limited transmission resources, and inter-user channel contention emerge as primary bottlenecks. The paper [17] proposes the end-to-end (E2E) uplink DeepJSCC-NOMA network for wireless image transmission; the paper [19] leverages NOMA with quantization modules to enable intelligent multi-user detection (IMUD); the work [20] employs semantic differential superposition coding at the transmitter and interference cancellation at the receiver (SeDSIC) to support both homogeneous and heterogeneous transmission (i.e., semantic information and bit streams); the paper [21] extracts and decorrelates both basic and enhanced source features, then superposes them probabilistically for transmission, allowing receivers to decode semantic information commensurate with channel quality; the authors of [22] propose semantic adaptive feature extraction (SAFE) network and integrated NOMA for downlink transmission; the work [23] integrates rate splitting, i.e., dividing information into public (non-orthogonal transmission) and private (orthogonal transmission) components, with generative adversarial networks (GAN)-based interference mitigation. However, these methods are still insufficient in terms of interference optimization in NOMA scenarios and cannot be applied to dynamic channel conditions. Additionally, the authors of [24] proposes Swin-Transformer-based feature fusion schemes under degraded broadcast channels (DBC) to reduce transmission overhead via joint representations, though the high computational complexity limits scalability in multi-user scenarios. Novel semantic multiple-access schemes have also emerged: the paper [25] exploits DeepJSCC’s interference resilience to propose an orthogonal model division approach, while the work [26] introduces semantic feature multiple access for multi-user video transmission. The authors of [27] propose a semantic feature domain interference management scheme for multi-user interference channels. Despite integrating traditional multiple-access principles with neural network robustness, these solutions lack systematic strategies for interference elimination and resource optimization in multi-user non-orthogonal settings. This paper focuses on uplink multi-user non-orthogonal semantic transmission and designs a low-complexity deep learning framework that can adapt to different channel conditions and further optimize to reduce the impact of interference.
II-B Knowledge Distillation and Its Application in Semantic Communication
II-C Diffusion Model and Its Application in Semantic Communication
Conventional DeepJSCC frameworks typically aim to minimize pixel-level distortion, such as mean squared error (MSE), as their optimization objective, but they ignore the improvement of image perceptual quality, especially in poor channel environments. In recent years, generative models have achieved groundbreaking progress in computer vision, with diffusion models emerging as a superior alternative to GANs due to their capacity to produce high-quality and semantically coherent images. Within the domain of semantic communication, GANs have been employed to improve image transmission quality. For instance, the paper [34] models the entire DeepJSCC transmission pipeline as the forward process, leveraging a pretrained StyleGAN2 to solve the inverse problem and improve perceptual quality. In contrast, diffusion models offer distinct advantages in stability, resistance to mode collapse, and generative diversity. Accordingly, the authors of [35] introduce the diffusion-aided JSCC framework, utilizing a pretrained Stable Diffusion model for the image reconstruction with initial JSCC decoding and multimodal conditions (spatial, textual, and channel state information). The work [36] introduces DiffCom which leverages channel received signal as condition to guide stochastic posterior sampling. Similarly, the works of [37] and [38] deploy pretrained diffusion models at the receiver to boost perceptual quality. Furthermore, the work [39] proposes channel denoising diffusion models (CDDM), which leverage diffusion models to preprocess channel outputs. The paper [40] presents the semantics-guided diffusion DeepJSCC (SGD-JSCC) scheme, which integrates diffusion models with multi-modal semantic guidance (i.e., text and edge maps) to design a channel denoising model. Inspired by these advancements, this paper innovatively proposes a two-stage image reconstruction framework tailored for semantic NOMA scenarios, deploying pretrained diffusion models at the receiver to achieve superior perceptual quality.
In sum, We identify critical research gaps in these domains and provide a comprehensive comparison with our KDD-SemNOMA framework in Table I.
TABLE I: Comparison with representative multi-user semantic communication methods
Method Dataset(s) Channel(s) Link direction Interference Management Generative Model Metric(s)
KDD-SemNOMA CIFAR-10, FFHQ-256 AWGN, Rayleigh Uplink Knowledge Distillation Diffusion Model PSNR, FID
DeepJSCC-NOMA [17] CIFAR-10 AWGN Uplink Curriculum Learning No PSNR
SemCom[33] Europarl Rayleigh Uplink Knowledge Distillation No BLEU
NOMASC[19] CIFAR-10, MINIST, Europarl AWGN, Rayleigh Downlink IMUD No PSNR, BLEU
IS-SNOMA[20] Cityscapes Rician Downlink SeDSIC No PSNR
DeepSCM[21] CIFAR-100 AGWN Downlink LMMSE Decorrelator No PSNR
DBC-Aware SC[24] CIFAR-10, STL-10, CelebA AWGN Downlink Semantic Fusion No PSNR
SAFE[22] ImageNet AWGN, Rayleigh Downlink Three-stage Training No PSNR
SFMA[26] CIFAR-10, SNU-FILM AWGN Downlink User Pairing Algorithm TAIN PSNR, LPIPS
SFDMA[27] MNIST, CelebA Rayleigh Bidirectional Information Bottleneck No PSNR
DeepPASIC[23] ImageNet AWGN Bidirectional Two-stage Transmission GAN PSNR
III System Model
III-A Overview
@@
The proposed KDD-SemNOMA framework is illustrated in Figure 1. Consider an uplink multi-user image semantic communication scenario where multiple user equipment (UE) transmit uplink image semantic information to the base station (BS). Unlike conventional communication systems employing separate source and channel coding, our framework adopts JSCC to directly transmit semantic features. Both UE and BS are configured with single antennas. Each UE transmits an image $\mathbf{x}_{i}\in\mathbb{R}^{C\times H\times W}$, where $i$ denotes the index of UE, $C$, $H$, and $W$ represent the color channels, width, and height of the RGB image, respectively. We denote $m=C\times H\times W$ as the total number of pixels in input image. The transmitter encodes the source image $\mathbf{x}_{i}$ into compressed semantic features $\mathbf{s}_{i}\in\mathbb{C}^{k}$ through a JSCC encoder, which can be expressed as
@@
Our framework for multi-user image reconstruction at the receiver comprises two primary stages: (1) multi-user DeepJSCC decoding optimized by knowledge distillation as detailed in Section IV, and (2) image refinement guided by diffusion-based generative priors as detailed in Section V. At the first stage, the SemNOMA model is trained using knowledge distillation to effectively decode transmitted signals. At the second stage, we leverage pretrained diffusion model priors to refine low-quality received images, where their visual fidelity and detail have further enhanced.
III-B Enhanced AF-Module
Wireless channels degrade semantic features via additive noise (AWGN channel) and multiplicative fading (Rayleigh channel), which distort feature distributions and impair reconstruction. In NOMA, these effects compound with multi-user interference, severely challenging feature recovery.
@@ III-B Enhanced AF-Module
Let the input semantic feature map be $\mathbf{f}_{s}\in\mathbb{R}^{B\times C^{\prime}\times H^{\prime}\times W^{\prime}}$, where $B$, $C^{\prime}$, $H^{\prime}$, and $W^{\prime}$ denote batch size, number of channels, height, and width, respectively. The channel side information is represented by $\mathbf{s}\in\mathbb{R}^{B\times D}$, with $D=3$ corresponding to the SNR $\gamma\in[0,20]$ dB and Rayleigh fading parameters of the $i$-th UE (amplitude $a_{i}$ and phase $\phi_{i}$). During training, $\gamma$ is uniformly sampled from $[0,20]$ dB, and the Rayleigh parameters $(a_{i},\phi_{i})$ are randomly generated to simulate diverse channel conditions.
@@
In summary, the enhanced AF-Module leverages channel state information to dynamically recalibrate features. It produces channel-aware attention masks that selectively emphasize robust components while suppressing noise-corrupted ones, enabling adaptive compensation for channel distortions.
III-C SemNOMA Network Architecture
@@
The encoder processes RGB images and normalizes them to the $[0,1]$ range by dividing pixel values by 255. It employs a stem layer for initial downsampling, followed by LayerNorm and additional downsampling stages. Multiple ConvNeXt blocks, with $N_{e}$ blocks per stage, process features, as detailed in Table II for $\times32$ resolution and Table III for $\times256$ resolution. Enhanced AF-Module, as introduced in Section III-B, improve the robustness to channel noise by conditioning on SNR and Rayleigh fading parameters (i.e., $\gamma,a,\phi$). The final $1\times 1$ convolution produces $M$ channels, normalized to meet the power constraint (i.e., Eq. 2), and reshaped into $k$ complex-valued channel inputs.
The decoder reconstructs image from the received multi-user superimposed semantic features. The $k$ complex-valued samples are converted to $2k$ real-valued channels, processed by a $1\times 1$ convolution and LayerNorm. Progressive upsampling stages then restore spatial resolution (e.g., $32\times 32$ or $256\times 256$), each followed by $N_{d}$ ConvNeXt blocks with parametric rectified linear unit (PReLU) activations and AF-Modules for feature refinement, as specified in Table II and Table III. A final $1\times 1$ convolution ($K=3$) with sigmoid activation produces a $C=3$ output, which is then denormalized to range $[0,255]$.
@@
IV Knowledge Distillation Optimization for Semantic NOMA (KD-SemNOMA)
@@
1:Multi-user image training dataset $\mathbf{x_{train}}=\{\mathbf{x}_{1}^{train},\dots,\mathbf{x}_{N}^{train}\}$, test dataset $\mathbf{x_{test}}=\{\mathbf{x}_{1}^{test},\dots,\mathbf{x}_{N}^{test}\}$, hyperparameters $\lambda_{1},\lambda_{2},\lambda_{3}$
@@
11:for each batch $\mathbf{x}_{\text{batch}}\subseteq\mathbf{x_{train}}$ do
@@
IV-A Teacher Model Training